Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework #19. §9.1 #1, 3, 5, 7, 9, 17, 19, 23, 25, 26.

• Homework Z (the homework formerly known as J). Due Wednesday, 17 April 2002. §9.1 #28.

• In section 9.2, we will only discuss tangents to parametric curves. We will not discuss area from this section.

• Homework #20. §9.2 #1, 3, 5, 6, 17, 23, 37, §8.2 #5, 17, 31 (I think you are to use your calculator to evaluate the integral in #31.) §9.3 #1, 3, 5, 7, 13, 15 (It might be fun to try to evaluate the integral in #15, but I hope you do not succeed.)

• Thursday, 18 April, The last quiz. §9.2.

• Homework #21. §9.4 #1, 3, 5, 7, 9, 11, 13, 15, 19, 25, 31, 33, 37, 39, 41, 51, 53, 78.

• In §9.4, you should familiarize yourself with polar coordinates, be able to convert to and from polar coordinates and practice plotting interesting curves in polar coordinates. We will not consider tangents to curves in polar coordinates.

• Homework #22 (the end). §9.5 #1, 3, 5, 7, 15, 17, 19, 21, 23, 27.

• In §9.5, we will find areas of regions represented in polar coordinates. A typical exam question might be something like #19 where you are to find the area enclosed by one loop of a curve.

• Homework 21 will be discussed on Thursday, 18 April 2002. Homework 22 will be discussed on Tuesday, 23 April 2002.

• Our final exam will be on Thursday, 2 May 2002 from 8am-10am in our regular classroom, CB234. The final exam will not be in Chemistry-Physics.

The final exam will consist of questions from the three midterms, and new material from Chapter 9 as listed on the syllabus and further clarified by the last few numbered homework assignments.

I will entertain questions about the proportion of new to old material at the end of the semester.

April 14, 2002